OpenStudy (anonymous):

Multiply and simplify: sqrt6(sqrt3+5sqrt2)

OpenStudy (bahrom7893):

3sqrt(2)+10sqrt(3)

OpenStudy (campbell_st):

expand the brackets \[\sqrt{6} \times \sqrt{3} = \sqrt{6\times3} = \sqrt{9 \times 2} = 3\sqrt{2}\] the other term is \[\sqrt{6} \times 5\sqrt{2} = 5\sqrt{6\times2} = 5\sqrt{4\times3} = 10\sqrt{3}\] answer is \[3\sqrt{2} + 10\sqrt{3}\]